OpenStudy (anonymous):

a baker has modeled the monthly operating costs for making wedding cakes by the function y=0.5x^2-12+150 where y is the total cost in dollars and x is the number of cakes prepared. what is the minimum cost?

OpenStudy (anonymous):

find the minimum cost by finding the second coordinate of the vertex the first coordinate is always \(-\frac{b}{2a}\) which in your case is \[-\frac{-12}{2\times .5}=12\]

OpenStudy (anonymous):

to find the actual cost, replace \(x\) by \(12\) and see what you get for the minimum. they really stretch for these "word" problems, don't they?